How do you solve 3x^2 - 6x = 2 by completing the square?
Answer by mananth(16946)

3x^2 - 6x = 2
/3
x^2 - 2x = 2/3
x^2-2x+1= 2/3 +1
(x-1)^2=5/3
take square root
x-1=+/-sqrt(5/3)
x= 1+sqrt(5/3)
OR
x=1-sqrt(5/3)
...
m.ananth@hotmail.ca